# HuggingGraph: Understanding the Supply Chain of LLM Ecosystem

### v1

@@ Abstract.
Large language models (LLMs) leverage deep learning architectures to process and predict sequences of words based on context, enabling them to perform a wide range of natural language processing tasks, such as translation, summarization, question answering, and content generation. However, the increasing size and complexity of developing, training, and deploying cutting-edge LLMs demand extensive computational resources and large-scale datasets. This creates a significant barrier for researchers and practitioners. Because of that, platforms that host models and datasets have gained widespread popularity. For example, on one of the most popular platforms, i.e., Hugging Face, there are more than 1.8 million models and more than 450K datasets by the end of Jun. 2025, and the trend does not show any slowdown.
As existing LLMs are often built from base models or other pre-trained models and using external datasets, they can inevitably inherit vulnerabilities, biases, or malicious components that exist in previous models or datasets. Therefore, it is critical to understand these components’ origin and development process to detect potential risks better, improve model fairness, and ensure compliance with regulatory frameworks. Motivated by that, this project aims to study such relationships between models and datasets, which are the central parts of the LLM supply chain. First, we design a methodology to collect LLMs’ supply chain information systematically. With the collected information, we design a new graph to model the relationships between models and datasets, which is a large directed heterogeneous graph, with 397,376 nodes and 453,469 edges.
Then, on top of this graph, we perform different types of analysis and make multiple interesting findings, such as (i) the LLM supply chain graph is large, sparse, and exhibits a power-law degree distribution; (ii) it features a densely connected core and a fragmented periphery; (iii) datasets are pivotal, playing critical roles in training; (iv) a strong interdependence exists between models and datasets; (v) the graph is dynamic, with daily updates capturing the ecosystem’s continuous evolution.
@@ 1. Introduction
Large language models (LLMs) are AI models designed to understand and generate human language by learning patterns and relationships within extensive datasets (Mirchandani et al., 2023; Shen et al., 2023), such as GPT (Generative Pre-trained Transformer) (Yenduri et al., 2024), BERT (Bidirectional Encoder Representations from Transformers) (Kenton and Toutanova, 2019), and T5 (Text-To-Text Transfer Transformer) (Bahani et al., 2023). These models leverage deep learning architectures to process and predict sequences of words based on context, enabling them to perform a wide range of natural language processing tasks (Bonner et al., 2023) such as translation (Lu et al., 2024b), summarization (Laban et al., 2023), question-answering (Allemang and Sequeda, 2024), and content generation (Agossah et al., 2023). LLMs usually have billions (or even trillions) of parameters (Merrick et al., 2024), enabling them to generate high-quality text.
@@
However, the increasing size and complexity of developing, training, and deploying cutting-edge LLMs demand extensive computational resources (Xu et al., 2024b) and large-scale datasets (Villalobos et al., 2024). This creates a significant barrier for researchers and practitioners, limiting their access to state-of-the-art models (Mizrahi et al., 2024). As the demand for democratizing access to such LLM models continues to rise, platforms that host models and datasets have gained widespread popularity. For example, Figure 1 shows the number of models and datasets (in a million scales) on Hugging Face, a popular AI model hosting platform (Face, [n. d.]), starting from July 31st, 2024, to Jun. 30th, 2025. By the end of June 2025, it has reached over 1.8M models and 450K datasets. In addition, the trend does not show any slowdown. Such LLM hosting platforms provide user-friendly interfaces, APIs, and cloud-based infrastructures that enable researchers and developers to easily share, fine-tune, and deploy models without requiring extensive computational resources. Moreover, they foster open collaboration, allowing the broader community to contribute to model improvements, benchmark performances, and enhance transparency in AI research.
@@
On such platforms, different types of models can be classified into two categories based on their tasks, i.e., base models and task-specific models. (i) base models are large, pre-trained models that can be fine-tuned for specific downstream tasks (Tan et al., 2024). They are usually trained on vast datasets and are general-purpose, such as GPT (Yenduri et al., 2024), BERT (Kenton and Toutanova, 2019), and T5 (Bahani et al., 2023). (ii) Task-specific models are modified versions of base models for a specific task. Taking Hugging Face as an example, there are four types of such models. First, the fine-tuned models adapt base models for specific tasks by training on additional task-specific datasets (Zou et al., 2023). Second, the adapters models add lightweight and modular layers to the pre-trained models for specific tasks (Hu et al., 2023). Third, the quantization models trade off the precision in numerical computations for accelerating inference and reducing memory consumption (e.g., using less precise model parameters) (Velingker et al., [n. d.]). Fourth, the merge models integrate multiple specialized or fine‑tuned parameter sets into a single unified model by combining weights or configurations (e.g., via weighted averaging or parameter stitching), enabling support for multiple tasks or domains without separate deployments (Akiba et al., 2025). Besides models, such platforms also host many datasets used for training and fine-tuning the previously discussed models (Röttger et al., 2024).
Figure 1. The number of models and datasets (in millions) on Hugging Face from July 31, 2024, to June 30, 2025.
@@ 1.1. Motivation
As existing LLMs are often built from base models or other pre-trained models and using external datasets, they can inevitably inherit vulnerabilities, biases, or malicious components that exist in previous models or datasets. Therefore, understanding these components’ origin and development process can help better detect potential risks, improve model fairness, and ensure compliance with regulatory frameworks.
Motivated by that, this project aims to study such relationships between models and datasets. They are the central parts of the LLM supply chain (Wang et al., 2024), which refers to the entire lifecycle of developing, training, and deploying LLMs, similar to a traditional supply chain in manufacturing or software development (Čolaković et al., 2021; Anonymous, 2022; Chou et al., 2006; Sonatype, 2015). The LLM supply chain can help to identify critical insights for both model evolution and dataset origin, as discussed below.
Model evolution. The study of the supply chain of LLMs gives a clear overview of how LLMs evolve from base models to fine-tuned variants, adapter integration, quantization models, and merged models that combine multiple specialized parameter sets into a single release. With that, one can easily keep track of them. For example, a use case is when a security or bias risk is found in one LLM, and we can quickly locate the potential models that might have the same problems.
Dataset origin. This supply chain can help to understand the datasets’ origins used for training different models (Shen et al., 2023). Dataset origin refers to the source from which the data is collected. For example, for a fine-tuned model, we not only care about which dataset is used for fine-tuning but also what other datasets are involved in training the previous model. Understanding such dataset origin helps to ensure that the dataset used is reliable, ethical, legally compliant, and relevant to the task.
@@ 1.2. Contribution
Our contributions are mainly threefold. First, we design a methodology to systematically collect the supply chain information of LLMs. In this paper, we mainly study the most popular ML platform, i.e., Hugging Face, but the same strategy also applies to other LLM platforms. In particular, we use the APIs from the ML platform to collect the metadata about the hosted model and dataset. To that end, we collected a large dataset, covering all the models (1.8M) and datasets (450K) as of June 30, 2025.
Second, with the collected metadata, we designed a new graph, named LLM supply chain graph, to model the relationships between models and datasets. It is a directed heterogeneous graph where a node denotes different types of datasets and models (including base, fine-tune, adapter, quantization, and merge models). An edge denotes the dependency relationship between them, including dataset-dataset, and model-dataset relationships. Together, this graph is able to accurately capture the LLM supply chain information. To this end, we constructed a large graph with 397,376 nodes and 453,469 edges. An anonymized version of the complete graph is publicly available at GitHub [^1] Anonymized LLM supply chain graph: https://github.com/huggingface00/HuggingGraph.
Third, with this graph, we perform different types of analysis, including model-dataset relationship analysis. We answer five research questions, including (i) the properties of the LLM supply chain graph, (ii) the structural analysis, (iii) the supply chain relationships between datasets, (iv) the supply chain relationships between the models and datasets, and (v) $\Delta$-based temporal update and tracking.
@@ 2. Preliminary
Note that this study mainly focuses on the relationships between various types of models and datasets, which are the central parts of the whole LLM supply chain ecosystem. We hope this study can not only provide insights on the relationships between LLM models and datasets but also raise awareness and future research interests in the LLM supply chain ecosystem.
@@ 3. Methodology
This section discusses the main methodologies, including data collection, LLM supply chain graph construction, and analysis.
@@
Model hub API Access the model hub to list, search, and download models and metadata.
Dataset API Access the datasets for discovery, metadata retrieval, and downloading.
Metrics API Access metrics for model evaluation, e.g., metric discovery, metadata retrieval, and calculation.
Search API Search the name, tag, or other metadata for models and datasets.
3.1. LLM Supply Chain Information Collection
Table 1 summarizes four types of APIs we used. In particular, (i) the model hub APIs allow one to access the hub of existing models, including searching and downloading the model and its metadata. (ii) The dataset APIs allow one to access the datasets for discovery, metadata retrieval, and downloading. (iii) The metrics APIs allow one to access the metrics for model evaluation, including metric discovery, metadata retrieval, and calculation. (iv) In addition, one can use search APIs to search the name, tag, or other metadata for models and datasets.
With the help of these APIs, we are able to collect the metadata of the hosted models and datasets. In particular, we collected a large dataset, covering the models (330K) and datasets (66K) as of June 30, 2025.
Handling missing information. Accurate construction of the LLM supply chain graph critically depends on the quality of metadata from the LLM platforms (e.g., Hugging Face), which might suffer from missing or incomplete data. To address this limitation, we design the following two techniques, i.e., cross-reference links and textual pattern extraction.
@@ 3.1. LLM Supply Chain Information Collection
Figure 2. An example subgraph centering on the base model “Meta-llama” from the complete LLM supply chain graph.
(ii) Textual pattern extraction. When structured metadata is absent, the model and dataset cards might mention dependencies in unstructured text descriptions. To capture that, we employ a named entity recognition (NER) method (Vasiliev, 2020; Li et al., 2020) to extract the dependency relationships from the text. For example, the textual phrase like “fine-tuned from Llama-2” contains the fine-tuned keyword, which implies from which model this model is actually fine-tuned. Similarly, we also look into other words, such as “train”, and “adapt”.
@@ 3.2. LLM Supply Chain Graph
With the collected metadata, we construct a directed heterogeneous graph to model the LLM supply chain. In this graph, a node denotes different types of datasets and models, including base, fine-tuned, adapter, and quantization models.
An edge denotes the dependency relationship between them, including model-model, dataset-dataset, and model-dataset relationships. Figure 2 shows a simplified supply chain subgraph centering on the base model Meta-llama.
(i) Dataset-dataset relationship. The datasets within the LLM supply chain might overlap, build upon, or extend from each other–e.g., a dataset may be a subset or modified version of another. To capture such information, we connect them with two types of edges. (1) Subset relationship. These arise when a dataset is explicitly documented as a subsample or partition of another. For instance, a dataset named “C4_200M” is described as a subset of “C4”. (2) Modified version. These represent updates or enhanced variants of existing datasets. For example, “TruthfulQA_v2” incorporates factual corrections and improvements over an earlier version, “TruthfulQA_v1”.
(ii) Model-dataset relationship. To capture the model and dataset dependency, we use the metadata information from both models and datasets. From the metadata of a model, it might include the datasets used to train it. However, not all the models disclose such information. To find more information, we observe the metadata of a dataset, including a data field on “trained_fine_tune_models” on this dataset. Thanks to that, we are able to capture the accurate model and dataset relationship. In Figure 2, the datasets The Pile and Awesome-Chatgpt-prompt have directed edges to model Meta-Llama, meaning that both datasets are used to train the model.
@@ 3.3. Supply Chain Graph Analysis
This supply chain graph can help to understand the transformational processes of the models and datasets. In particular, we can understand how base models evolve into their variants, including fine-tuned, adaptive, and quantized models, and vice versa. Similar observations can be made for datasets. This would provide a clear view of how the base model (or dataset) is transformed for performing a particular task. In particular, we mainly perform two types of analysis, i.e., forward and backward analysis.
@@
Dataset analysis example. For the dataset, our supply chain analysis shows how different datasets connect and form a new dataset and work together. This combination creates flexible resources that improve how models perform in various areas. In Figure 2, the dataset The Pile is composed of multiple subsets, including Wikimedia, Arxiv, Openwebtext2, and Pubmed Central. Together, these datasets form a unified corpus that serves as training data for models like Meta-LlaMA.
@@
Model and dataset analysis example. In Figure 2, to analyze the backward supply chain of the model RBot70Bv4, we trace its lineage back through its development stages. This model is fine-tuned from Unsloth, which in turn originates from its base model, Meta-Llama, and the datasets used to train the base model are The Pile and Awesome-Chatgpt-prompts. Through this analysis, we establish the backward path, starting from the target model—RBot70Bv4—and tracing back to its base model, Meta-Llama, revealing the dependencies and transformations involved in its development. Similarly, we can analyze the datasets.
3.4. Temporal Update and Tracking
To ensure HuggingGraph remains an accurate and evolving representation of the LLM ecosystem, we have implemented a (i) $\Delta$-based batch update mechanism and (ii) a versioned archival system.
$\Delta$-based graph update. To ensure HuggingGraph reflects the evolving nature of the LLM ecosystem, we implement a $\Delta$-based batch update mechanism. Instead of reconstructing the graph from scratch, we compute incremental updates by comparing successive versions. Let $G_{t}$ denote the graph at time $t$, and define the $\Delta$ as $\Delta_{t}=G_{t}-G_{t-1}$, where $\Delta_{t}$ captures all added or removed nodes and edges. This approach offers both computational scalability and storage efficiency. At scheduled intervals (e.g., daily), we collect metadata from Hugging Face and compute the differences. For instance, between June 25–July 15, 2025, the platform added 104,380 new models ($\sim$3,866/day), 20,034 new datasets ($\sim$742/day), and approximately 854 dependency edges/day.
Versioning and longitudinal tracking. To enable temporal and historical analyses, each update generates a versioned snapshot of the graph, archived with a timestamp and associated change log. This log captures additions, deletions, and dependency modifications, such as new edges from TruthfulQA v2 to TruthfulQA v1. This ensures that HuggingGraph remains a living resource that adapts to ongoing developments in LLM research and deployment.
4. Experiment and Finding

### v2

@@ Abstract.
Large language models (LLMs) leverage deep learning architectures to process and predict sequences of words based on context, enabling them to perform a wide range of natural language processing tasks, such as translation, summarization, question answering, and content generation. However, the increasing size and complexity of developing, training, and deploying cutting-edge LLMs demand extensive computational resources and large-scale datasets. This creates a significant barrier for researchers and practitioners. Because of that, platforms that host models and datasets have gained widespread popularity. For example, on one of the most popular platforms, i.e., Hugging Face, there are more than 1.8 million models and more than 450K datasets by the end of June 2025, and the trend does not show any slowdown.
As existing LLMs are often built from base models or other pre-trained models and use external datasets, they can inevitably inherit vulnerabilities, biases, or malicious components that exist in previous models or datasets. Therefore, it is critical to understand these components’ origin and development process to detect potential risks better, improve model fairness, and ensure compliance with regulatory frameworks. Motivated by that, this project aims to study such relationships between models and datasets, which are the central parts of the LLM supply chain. First, we design a methodology to collect LLMs’ supply chain information systematically. With the collected information, we design a new graph to model the relationships between models and datasets, which is a large directed heterogeneous graph, having 402,654 nodes and 462,524 edges. Then, on top of this graph, we perform different types of analysis and make multiple interesting findings.
@@ 1. Introduction
Large language models (LLMs) are AI models designed to understand and generate human language by learning patterns and relationships within extensive datasets (Mirchandani et al., 2023; Shen et al., 2023), such as GPT (Generative Pre-trained Transformer) (Yenduri et al., 2024), BERT (Bidirectional Encoder Representations from Transformers) (Kenton and Toutanova, 2019), and T5 (Text-To-Text Transfer Transformer) (Bahani et al., 2023). These models leverage deep learning architectures to process and predict sequences of words based on context, enabling them to perform a wide range of natural language processing tasks (Bonner et al., 2023), such as, translation (Lu et al., 2024), summarization (Laban et al., 2023), question-answering (Allemang and Sequeda, 2024), and content generation (Agossah et al., 2023). LLMs usually have billions (or even trillions) of parameters (Merrick et al., 2024), enabling them to generate high-quality text.
@@
However, the increasing size and complexity of developing, training, and deploying cutting-edge LLMs demand extensive computational resources (Xu et al., 2024b) and large-scale datasets (Villalobos et al., 2024). This creates a significant barrier for researchers and practitioners, limiting their access to state-of-the-art models (Mizrahi et al., 2024). As the demand for democratizing access to such LLM models continues to rise, platforms that host models and datasets have gained widespread popularity. For example, Figure 1 shows the number of models and datasets (in a million scales) on Hugging Face, the largest public AI model hosting platform (Face, [n. d.]), starting from July 2024 to June 2025. By the end of June 2025, it has reached over 1.8M models and 450K datasets. In addition, the trend does not show any slowdown. Such LLM hosting platforms provide user-friendly interfaces, APIs, and cloud-based infrastructures that enable researchers and developers to easily share, fine-tune, and deploy models without requiring extensive computational resources. Moreover, they foster open collaboration, allowing the broader community to contribute to model improvements, benchmark performances, and enhance transparency in AI research.
@@
On such platforms, different types of models can be classified into two categories based on their tasks, i.e., base models and task-specific models. (i) Base models are large, pre-trained models that can be fine-tuned for specific downstream tasks (Tan et al., 2024). They are usually trained on vast datasets and are general-purpose, such as GPT (Yenduri et al., 2024), BERT (Kenton and Toutanova, 2019), and T5 (Bahani et al., 2023). (ii) Task-specific models are modified versions of base models for a specific task. Taking Hugging Face as an example, there are four types of such models. First, fine-tuned models adapt base models for specific tasks by training on additional task-specific datasets (Zou et al., 2023). Second, adapters models add lightweight and modular layers to the pre-trained models for specific tasks (Hu et al., 2023). Third, quantization models trade off precision in numerical computations for accelerating inference and reducing memory consumption (e.g., using less precise model parameters) (Velingker et al., [n. d.]). Fourth, merged models integrate multiple models into a single unified model by combining weights or configurations enabling support for multiple tasks or domains without separate deployments (Akiba et al., 2025). Besides models, such platforms also host many datasets used for training and fine-tuning the previously discussed models (Röttger et al., 2024).
@@ 1.1. Motivation
As existing LLMs are often built from base models or other pre-trained models and using external datasets, they can inevitably inherit vulnerabilities, biases, or malicious components from previous models or datasets. Thus, understanding these components’ origin and provenance can help better detect potential risks, improve model fairness, and ensure compliance with regulatory frameworks.
Figure 1. The number of AI models and datasets (in a million scale) on Hugging Face from July 2024 to June 2025.
Motivated by that, this paper aims to study such relationships between models and datasets. They are the central parts of the LLM supply chain (Wang et al., 2024), which refers to the entire lifecycle of developing, training, and deploying LLMs, similar to a traditional supply chain in manufacturing or software development (Čolaković et al., 2021; Anonymous, 2022; Chou et al., 2006; Sonatype, 2015). Such a supply chain can help to identify critical insights for both model evolution and dataset origin, as discussed below.
Model evolution. The study of the LLM supply chain gives a clear overview of how LLMs evolve from base models to fine-tuned variants, adapter integration, and quantization models. With that, one can easily keep track of them. For example, a use case is when a security vulnerability is found in one LLM, and we can quickly locate the potential models that might have the same issues.
Dataset origin. This supply chain can also help to understand the datasets’ origins used for training different models (Shen et al., 2023). Dataset origin refers to the source from which the data is collected. For example, for a fine-tuned model, we not only care about which dataset is used for fine-tuning but also what other datasets are involved in training the previous model. Understanding such dataset origin helps to ensure that the dataset used is reliable, and legally compliant.
@@ 1.2. Contribution
Our main contributions are threefold. First, we design a methodology to systematically collect the supply chain information of LLMs. In this paper, we mainly study the most popular AI platform, i.e., Hugging Face, but the same strategy applies to other platforms. In particular, we use the APIs from the AI platform to collect the metadata about the hosted model and dataset. To this end, we collected a large dataset as of June 30, 2025.
Second, with the collected metadata, we created a new graph, named LLM supply chain graph, to model the relationships between models and datasets. It is a directed heterogeneous graph where a node denotes different types of datasets and models (including base, fine-tune, adapter, quantization, and merge). An edge denotes the dependency relationship between them, including model-model, dataset-dataset, and model-dataset relationships. Together, this graph is able to accurately capture the LLM supply chain information. To this end, we constructed a large graph with 402,654 nodes and 462,524 edges. An anonymized version of the complete graph is publicly available at GitHub [^1] LLM supply chain graph: https://github.com/huggingface00/HuggingGraph.
Third, with this graph, we perform different types of analysis, including forward and backward analysis. We study seven research questions, including (i) the properties of the LLM supply chain graph, (ii) structural analysis, (iii) supply chain relationships between AI models, (iv) supply chain relationships between datasets, (v) supply chain relationships between models and datasets, (vi) dynamic update evaluation, and (vii) extension to other AI platforms.
@@ 2. Preliminary
This study mainly focuses on the relationships between models and datasets, which are the central parts of the whole LLM supply chain ecosystem. We hope this study can not only provide insights on the LLM supply chain but also raise awareness and future research interests in this direction.
@@ 3. Methodology
3.1. LLM Supply Chain Information Collection
Table 1 summarizes the four types of APIs we used. In particular, (i) the model hub APIs allow access to the hub of existing models, including searching and downloading the model and its metadata. (ii) The dataset APIs allow access to the datasets for discovery, metadata retrieval, and downloading. (iii) The metrics APIs allow access to the metrics for model evaluation, including metric discovery, metadata retrieval, and calculation. (iv) One can use the search APIs to search the name, tag, or other metadata for models and datasets.
@@ 3.1. LLM Supply Chain Information Collection
Model hub Access the model hub to list, search, and download models and metadata.
Dataset Access the datasets for discovery, metadata retrieval, and downloading.
Metrics Access metrics for model evaluation, e.g., metric discovery, metadata retrieval, and calculation.
Search Search name, tag, or other metadata for model and dataset.
Handling missing information. Accurate construction of the LLM supply chain graph critically depends on the quality of metadata from the LLM platforms (e.g., Hugging Face), which might suffer from missing or incomplete data. To address this limitation, we apply the following two techniques, i.e., cross-reference links, and textual pattern extraction.
@@
(ii) Textual pattern extraction. When structured metadata is absent, the model and dataset cards might mention dependencies in unstructured text descriptions. To capture that, we employ a named entity recognition (NER) method (Vasiliev, 2020; Li et al., 2020) to extract the dependency relationships from the text. For example, the textual phrase like “fine-tuned from Llama-2” contains the fine-tuned keyword, which implies from which model this model is actually fine-tuned. Similarly, we also look into other words, such as, “train”, and “adapt”.
@@ 3.2. LLM Supply Chain Graph
With the collected metadata, we construct a directed heterogeneous graph to model the LLM supply chain. In this graph, a node denotes different types of datasets and models, including base, fine-tune, adapter, quantization, and merge models. An edge denotes the dependency relationship between them, including model-model, dataset-dataset, and model-dataset relationships.
Figure 2 shows a simplified supply chain subgraph centering on a base model Meta-llama. (i) Model-model relationship. To further identify the supply chain relationship, we will check the relevant data fields. In particular, given a model in Hugging Face, there are data fields “finetune” that show which models are fine-tuned from this model. Similarly, “adapter”, “quantization”, and “merge show which models are adapted, quantized or merged from them, respectively. With such information, we can construct the supply chain relationship between the models. As shown in Figure 2, model Llama-3.3-70B is fine-tuned from the base model Meta-llama. Then, it is used by the models Doctor-Shotgun and Llama-3.3-70B-4bit-Vision to generate an adapter and quantization model, respectively.
(ii) Dataset–dataset relationship. The datasets within the LLM supply chain might overlap, build upon, or extend from each other. For example, a dataset may be a subset or modified version of another. To capture such information, we connect them with two types of edges. (1) Subset relationships arise when a dataset is explicitly documented as a subsample or partition of another. For instance, a dataset named “C4_200M” is described as a subset of “C4”. (2) Modified versions represent updates or enhanced variants of existing datasets. For example, “TruthfulQA_v2” incorporates corrections and improvements over an earlier version, “TruthfulQA_v1”.
(iii) Model-dataset relationship. To capture the model and dataset dependency, we use the metadata from both models and datasets. The metadata of a model might specify the datasets used for training or adapting. However, not all the models disclose such information. To capture more information, we find the metadata of a dataset contains a data field of “trained_fine_tune_models” on this dataset. Thanks to that, we are able to capture the accurate model and dataset relationships. In Figure 2, the datasets The Pile and Chatgpt-prompt have directed edges to model Meta-Llama, meaning that both datasets are used to train the model.
Figure 2. An example subgraph centering on base model “Meta-llama” from the complete LLM supply chain graph.
@@ 3.3. Supply Chain Graph Analysis
This supply chain graph can help to understand the transformational processes of the models and datasets. In particular, we can understand how base models evolve into their variants, including fine-tuned, adaptive, quantized, or merged models, and vice versa. Similar observations can be made for datasets. This would provide a clear view of how the base model (or dataset) is transformed for performing a particular task. In particular, we mainly perform two types of analysis, i.e., forward and backward analysis.
Forward analysis is the method of traversing the supply chain graph following the dependency edges of a chosen node in a forward-going way. This node (known as the root/source node) can be a dataset, a base model, a fine-tuned model, an adapter, or a quantized model. In particular, given a source node, we apply the graph traversal algorithm (e.g., breadth-first search (BFS) (Tarjan, 1972)) to traverse all the nodes (including both models and datasets) in a level-by-level pattern. To that end, this forward analysis will identify all the nodes that are reachable from the source node.
Model analysis example. In Figure 2, to analyze the forward supply chain of Meta-Llama, we trace its development from training datasets to subsequent adaptations. This model is fine-tuned into Llama-3.3-70B, enhancing its capabilities for specific tasks. Summarizing the forward supply chain of Meta-Llama, we identify the following four distinct forward supply chains: (i) Base model (Meta-llama) $\rightarrow$ fine-tuned model (Llama-3.3-70B) $\rightarrow$ another fine-tuned model (Llama3.3-70B-Vision). (ii) Base model (Meta-llama) $\rightarrow$ adapted model (Doctor-Shotgun). (iii) Base model (Meta-llama) $\rightarrow$ fine-tuned model (Llama-3.3-70B) $\rightarrow$ quantization model (Llama-3.3-70B-4bit). (iv) Base model (Meta-llama)) $\rightarrow$ merged model (MistLlama). These pathways illustrate the evolution trajectory of the base model Meta-Llama showcasing its progressive specialization and adaptation for various tasks.
@@
Dataset analysis example. For the dataset, our supply chain analysis shows how different datasets connect and form a new dataset. This combination creates flexible resources that show how models perform in various areas. In Figure 2, the dataset The Pile is composed of multiple subsets, including Wikimedia, Arxiv, Openwebtext2, and Pubmed Central. Together, these datasets form a unified corpus that serves as training data for models like Meta-LlaMA.
Backward analysis, on the other hand, is the method of traversing the supply chain graph following the edges in a backward way. We accomplish this by traversing the directed graph also with BFS (Knuth, 1974) in reverse, starting from the selected node and following the incoming edges. To that end, this backward analysis will identify all the nodes that can reach the source node.
@@
Model and dataset analysis example. In Figure 2, analyzing the backward supply chain of model RBot70Bv4, we trace its lineage through its development stages. This model is fine-tuned from Unsloth, which in turn originates from its base model, Meta-Llama, and the datasets used to train the base model are The Pile and Chatgpt-prompts. Through this analysis, we establish the backward path, starting from the target model, i.e., RBot70Bv4, and tracing to its base model, Meta-Llama, revealing dependencies and transformations involved in its development. Similarly, we can analyze datasets.
3.4. Accommodating Dynamic Update
The hosted models and datasets on AI platforms are growing fast as new models are being developed every day. For example, between June 25 and July 15, 2025, we observed 80,703 new models (approximately 3,843 per day), 27,405 new datasets (approximately 1,305 per day) on Hugging Face only, which is just one of the many AI platforms. Therefore, we need to accommodate the dynamic update to accurately manage and analyze the AI supply chain.
Particularly, HuggingGraph accommodates the dynamic update in three steps. (i) Scoping updated models or datasets. At a time $t$, we keep a copy of the hosted models and datasets with their IDs. When we evolved to $t+1$, we get another copy of the hosted models and datasets with their IDs. The difference between them shows the updated models and datasets, including newly added or deleted. In our current implementation, we are keeping the update on a daily basis. (ii) Metadata collection for the updated models or datasets. For the identified updated models or datasets, we will collect their metadata using the same strategy as discussed in Section 3.1. To this end, we get the updated dependencies between models and datasets. That is, for the update at time $t+1$ compared to $t$, it can be represented as $\Delta_{t+1}$. (iii) $\Delta$-based dynamic graph update. Given the newly updated dependency $\Delta_{t+1}$, and let $G_{t}$, $G_{t+1}$ denote the graph at time $t$, $t+1$, respectively, then $G_{t+1}=G_{t}\cup\Delta_{t+1}$.
4. Experiments and Finding